OpenStudy (kewlgeek555):

URGENT - Please Help?! Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (kewlgeek555):

a. y = four thirdsx + 8 b. y = four thirdsx + 3 c. y = -three fourthsx + 8 d. y = -three fourthsx + 3

OpenStudy (kewlgeek555):

Sorry for the whole "URGENT" thing - I need this ASAP.

OpenStudy (kewlgeek555):

I actually tried to figure this out - I can show you my work if you'd like.

OpenStudy (anonymous):

Your answer is a The gradient is 4/3, you can see that if you write it in the form y=mx + c where m is the gradient. So y = 4/3 (x) + 12 Plug in -3 you get y to be 8 :)

OpenStudy (kewlgeek555):

Oh so you PLUG IN -3?

OpenStudy (anonymous):

Yeah it has to pass through that point, and the gradient has to be the same, so it will cross the axis at a different point, you could put -3 or 4 in, you'll get the same answer

OpenStudy (kewlgeek555):

Oh okay. Thank you so much. Do you think you could help me on couple more?

OpenStudy (anonymous):

Sure

OpenStudy (kewlgeek555):

Thanks. http://orange.flvs.net/webdav/assessment_images/educator_algebra1_v17/03_07_quiz/03_07_14.jpg Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). a. y = 2x + 4 b. y = - 2x + 2 c. y = - 1 over 2 x + 2 d. y = 2x - 1 Now, I know that it cannot be B or C because for this new line to be perpendicular, it had to have a positive slope, right? I think it would be reasonable to say D, am I correct?

OpenStudy (anonymous):

You are right, it has to have a positive slope, however I don't think it's D, If it has to pass through (-1,2) then D, which says y=2x -1 should satisfy these points So y = 2(-1) + 2 = 0, this does not equal 2 However a, y = 2x +4 y = 2(-1) + 4 = -2 + 4 = 2 Does satisfy your points. Do you see how much easier it just to plug in your values and see if they work? :)

OpenStudy (kewlgeek555):

So it would be A? Okay! I have another one. (SORRY!) Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3

OpenStudy (anonymous):

It's c

OpenStudy (anonymous):

Can you think why? Give me an explanation!

OpenStudy (kewlgeek555):

Okay. I think it is because...ooh, no. Well I see that it was three halves and now it is 2/3s. I think you might have plugged in -6.

OpenStudy (anonymous):

If it's a line perpendicular to a line you want you always 'flip' the gradient and times by -1. This will make it intersect at 90 degrees :) So a perpendicular gradient to 3/2 is -2/3! A perpendicular gradient to 10/7 is -7/10 So it has to be one of the last two, now you can put it in your -6 for x, and see which one of those two gives you y=3, it is the third one :)

OpenStudy (kewlgeek555):

Okay. I have one of those questions I want to solve by myself. I am going to show you. Then, I have another one I am stuck on. Here is the one I worked on myself. Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x - 7 y = - one halfx - 2 y = 2x - 2 y = - one halfx - 7 "parallel to y = 2x +2" "passes through the point (5,3)" y = mx + b, where m = 2, so... y = 2x + b. I plug in 5 for b. 3 = 2(5) + b 3 = 10 + b -7 = b Answer y = 2x - 7

OpenStudy (anonymous):

Yeah, well done :)

OpenStudy (kewlgeek555):

:-D

OpenStudy (kewlgeek555):

Well, I also have one more question.

OpenStudy (anonymous):

Sure :)

OpenStudy (kewlgeek555):

http://orange.flvs.net/webdav/assessment_images/educator_algebra1_v17/03_07_quiz/03_07_16.jpg Select the equation of the line that passes through the point (-2, 3) and is perpendicular to the line on the graph. y = 0 y = -2 y = 3 y = 3x

OpenStudy (kewlgeek555):

That is the last one.

OpenStudy (anonymous):

Are you sure it's (-2,3) ?

OpenStudy (kewlgeek555):

Yep, I am.

OpenStudy (anonymous):

Ok, then the answer is c, y = 3

OpenStudy (anonymous):

Can you see how it couldn't be y = 3x, it wouldn't be perpendicular. And the only one of the other three which gives you a y coordinate of 3 is y=3

OpenStudy (kewlgeek555):

Okay. Thans. and I am going to write all the tips you gave me in my notes. ;]

OpenStudy (anonymous):

Glad I could help :)